# RESVMUNetX: A Low-Light Enhancement Network Based on VMamba

### v1

@@ 3 Method
$Step1:pred=estimaiton\ error+initial\ input$ (1)
$Step2:pred_{denoise}=estimaiton\ error+initial\ input_{denoise}$ (2)
@@ 3.1 VMUNetX
Figure 3: The overview of Residual Mixed Upsampling
Residual Mixed Upsampling: In the mixed upsampling module, as shown in Figure 3,a strategy that combines direct upsampling and error regression is adopted, where error regression utilizes a method of residual connections to effectively enhance the feature maps of low-light images. This module is capable of increasing the size of the feature map while simultaneously ensuring dense transfer of feature information and effective control of the number of channels. It features two parallel branches, one responsible for enhancing features through a convolutional network and performing upsampling, and the other directly achieving upsampling through bilinear interpolation. The outputs of the two branches are merged by addition at the end, aiming to combine information of two different characteristics, thereby enriching the expressive capability of the final upsampled feature map. This design concept not only draws on the residual learning approach from traditional deep learning to enhance the model’s learning capacity and accelerate training convergence but also utilizes the direct upsampling branch to preserve more original feature information. Through the mixed strategy, the upsampling module can more effectively retain important information in the image during the feature enhancement and upsampling process, making it suitable for low-light image enhancement domains that require high detail in images.
@@ 4.2 Low-light Image Enhancement
Table 1: Performance metrics compared for different methods on LOLV1 and LOLV2 datasets (real and synthetic).
@@ 4.3 Ablation Study
On the LOL v1 dataset, we conducted a series of ablation studies aimed at evaluating the performance of five different network architectures. These architectures were modified by removing or adding specific components to compare their respective image enhancement effects.
Table 3: Performance metrics compared for different VMUNet configurations.

### v2

@@ 3 Method
$Step1:pred=Error\ estimation+Low\ light\ input$ (1)
$Step2:pred_{denoise}=Error\ estimation+Denoising\ input$ (2)
@@ 3.1 VMUNetX
Residual Mixed Upsampling: In the mixed upsampling module,a strategy that combines direct upsampling and error regression is adopted, where error regression utilizes a method of residual connections to effectively enhance the feature maps of low-light images. This module is capable of increasing the size of the feature map while simultaneously ensuring dense transfer of feature information and effective control of the number of channels. It features two parallel branches, one responsible for enhancing features through a convolutional network and performing upsampling, and the other directly achieving upsampling through bilinear interpolation. The outputs of the two branches are merged by addition at the end, aiming to combine information of two different characteristics, thereby enriching the expressive capability of the final upsampled feature map. This design concept not only draws on the residual learning approach from traditional deep learning to enhance the model’s learning capacity and accelerate training convergence but also utilizes the direct upsampling branch to preserve more original feature information. Through the mixed strategy, the upsampling module can more effectively retain important information in the image during the feature enhancement and upsampling process, making it suitable for low-light image enhancement domains that require high detail in images.
@@ 4.2 Low-light Image Enhancement
Table 1: Performance metrics compared for different methods on LOLV1 and LOLV2 datasets (real and synthetic).The highest results are in red color.
@@
4.3 Low-light Object Detection
Table 3: Low-light detection results on ExDark enhanced by different algorithms.The highest results are in red color.
Method Bicycle Boat Bottle Bus Car Cat Chair Cup Dog Motorbike People Table All
SCI 65 67.5 54.3 77.3 70 56.8 56.2 52.6 46.6 47.7 70.2 50.4 59.5
Zero-Dce 63 65.9 52.7 75.2 69.9 52.9 56.1 50.4 50.4 43.5 69.9 50 58.3
Retinexformer 66.9 66.5 54.5 78.4 69.8 51.5 53.6 51.8 51.6 49.4 70.3 52.1 59.7
SMNet 62.2 66.9 52.1 75.4 70.8 51.9 55 56.2 50.3 49.3 70.5 49.7 59.2
RUAS 62.8 62.2 51.4 71.5 62.2 51.5 50.1 45.6 45.1 48.4 64.5 48.3 55.3
ResVMUNetX 61.6 65.8 55.5 74 70.3 57.7 55.3 53.2 52.4 48.2 72 53.7 60
Experiment Settings: We conducted low-light object detection experiments on the Exdark dataset to quantitatively and qualitatively compare the preprocessing effects of different low-light enhancement algorithms on advanced visual tasks. The Exdark dataset consists of a total of 7363 underexposed photos, with 12 object categories annotated with bounding boxes. In the experiments, 5884 images were used for training, and the remaining 1479 images were used for testing. For various low-light enhancement algorithms, we trained them on the LOLv1 dataset and used them as fixed-parameter preprocessing modules to preprocess the Exdark dataset. As for the detector, we used YOLOv8 and fixed the relevant training parameters before training.
Quantitative Results: The map50 scores of different enhancement algorithms are shown in Table 3. Our ResVMUNetX achieved the highest map50 score, with 60 AP. It outperformed the best self-supervised algorithm by 0.5 AP and the current best fully supervised algorithm by 0.3 AP. Additionally, ResVMUNetX performed the best on 5 object categories:Bottle,Cat,Dog,People and Table.
Figure 4: Visual comparison of object detection in low-light (left) and enhanced (right) scenes by our method on the Exdark dataset.
Qualitative Results: Figure 4 shows a visual comparison of detection results in low-light scenes and scenes enhanced by ResVMUNetX. In low-light scenes, the detector is prone to missing some objects due to underexposure, such as the bottle in (a) and the underexposed car in (b).And it will also cause misidentification of the object due to insufficient exposure, as shown in c. In contrast, after enhancement by ResVMUNetX, the detector can better capture objects in the images and reliably detect them. The anchor boxes are also more accurate, indicating that our method has a certain enhancement effect on high-level visual tasks.
@@ 4.4 Ablation Study
On the LOL v1 dataset, we conducted a series of ablation studies aimed at evaluating the performance of five different network architectures. These architectures were modified by removing or adding specific components to compare their respective image enhancement effects.The data from the relevant ablation experiments can be visualized in Table 4.
Table 4: Performance metrics compared for different VMUNet configurations.The highest results are in red color.